Calculate the latent heat of vaporization for water at normal boiling point and at 250oC using Riedel and Watson equations. Compare with values from steam table.
Answer #1

Calculation of heat of vapourization of water at normal boiling point :

Using Riedel's equation

H1 = 1.093×RT1 ( lnPc - 1.013 ) / { 0.930 - (T1/Tc) }

where, H1 = heat of vapourization of water at normal boiling point

R = gas constant

T1 = normal boiling point of water

Tc = critical temperature of water

Pc = critical pressure of water

Using the standard values of

R = 8.314 J/ K.mol ,

T1 = 373 K,

Tc = 647 K

and Pc = 221 bar , in the Riedel's equation

H1 = 1.093×8.314×373{ln221 - 1.013} / {0.930 - (373/647)}

H1 = 3389.526 {5.398 - 1.013} / {0.930 - 0.576}

H1 = 3389.526× 4.385 / 0.35422

H1 = 41986 J/mol

Let H2 = heat of vapourization of water at 250°C

T2 = 250°C = 523 K

Using watson equation,

H2 / H1 = [ (Tc - T2) / (Tc - T1 ) ]0.38

H2 / 41986 = [ (647 - 523) / (647 - 373) ]0.38

H2 / 41986 = [ 0.45255 ]0.38

H2 = 0.73987 × 41986

H2 = 31064.18 J/mol